OpenStudy (anonymous):

Which polynomial function has a leading coefficient of 1 and roots (7 + i) and (5 – i) with multiplicity 1?

OpenStudy (anonymous):

f(x) = (x + 7)(x – i)(x + 5)(x + i) f(x) = (x – 7)(x – i)(x – 5)(x + i) f(x) = (x – (7 – i))(x – (5 + i))(x – (7 + i))(x – (5 – i)) f(x) = (x + (7 – i))(x + (5 + i))(x + (7 + i))(x + (5 – i))

OpenStudy (anonymous):

please help will fan and double medal